OpenStudy (anonymous):

What is the local minimum of 4x^3+6x^2-24x+6

OpenStudy (anonymous):

and local maximum?

OpenStudy (anonymous):

y = = 4x^3 + 3x^2 - 6x + 1 y´ = 12x² + 6x - 6 = 0 6(2x² + x - 1) = 0 (2x - 1)(x + 1) = 0 x = 1/2, -1 y´´ = 24x + 6 y´´(-1) = 24(-1) + 6 = -18 negative, this is a maximum y´´(1/2) = 24(1/2) + 6 = 18 positive, this is a minimum y = 4(-1)^3 + 3(-1)^2 - 6(-1) + 1 = 6 y = 4(1/2)^3 + 3(1/2)^2 - 6(1/2) + 1 = -3/4 Maximum.....(-1, 6) <================== Minimum......(1/2, -3/4) <==================